OpenStudy (anonymous):

Help solve this equation for x (x/(x-2)+(2x/(4-x^2)=(5/(x+2)

OpenStudy (anonymous):

Clear denominators by mult. through by the LCD=(x+2)(x-2) to get\[x(x+2)-2x=5(x-2)\]You can then distribute and collect like terms to solve. (cont.)

OpenStudy (anonymous):

Then use quadratic formula to solve \[x^2-5x-10=0\](cont.)

OpenStudy (anonymous):

See attached for quadratic solution